# The GPT-WritingPrompts Dataset: A Comparative Analysis of Character Portrayal in Short Stories

### v1

@@ Abstract
The improved generative capabilities of large language models have made them a powerful tool for creative writing and storytelling. It is therefore important to quantitatively understand the nature of generated stories, and how they differ from human storytelling. We augment the Reddit WritingPrompts dataset with short stories generated by GPT-3.5, given the same prompts. We quantify and compare the emotional and descriptive features of storytelling from both generative processes, human and machine, along a set of six dimensions. We find that generated stories differ significantly from human stories along all six dimensions, and that human and machine generations display similar biases when grouped according to the narrative point-of-view and gender of the main protagonist. We release our dataset and code at https://github.com/KristinHuangg/gpt-writing-prompts.
@@ 1 Introduction
Storytelling is a fundamental human practice, and a rich representation of the beliefs, practices, and languages of communities. The stories of a particular era often reflect the social norms and biases of the time, and also function as a space for writers to imagine and explore alternative worlds. The characters [^1] We use the term ‘character’ in this work to refer to fictional people, i.e, named entities that speak and act in stories. in a story are often central to this process: protagonists are usually depicted with socially and personally desirable traits like kindness, intelligence, and strength, and villains with undesirable characteristics (greed, cruelty, apathy, weakness). The domain of literature therefore functions as a rich source of cultural information for researchers in the social sciences, ranging from historians to linguists Piper et al. (2021); Ziems et al. (2024). Consequently, we have also seen extensive research on NLP methods that can model and understand these aspects of stories, and draw inferences about social biases represented in them Fast et al. (2016b); Lucy and Bamman (2021); Sap et al. (2022).
Large language models (LLMs), pretrained on billions of tokens of text from the internet, are now able to generate several thousand words of coherent output in response to user-specified input contexts Bommasani et al. (2021). Naturally, these systems have become popular tools for a number of tasks involving text generation, both controlled and open-ended. e.g., dialogue systems, code generation, summarization, question answering, and creative writing. However, neural models have been shown to replicate and amplify the biases encoded in their training data Bender et al. (2021). It is therefore interesting – and important – to understand the nature of stories that are told by these models, and how they differ from human models of storytelling Chakrabarty et al. (2023); Xie et al. (2023).
@@
    Extending the Reddit WritingPrompts dataset Fan et al. (2018) to include generations from OpenAI gpt-3-turbo (among the later releases from OpenAI’s set of GPT models). The original dataset consists of 97,222 unique writing prompts and multiple stories written by Reddit users in response to each prompt. We generate a corresponding set of artificial stories by prompting gpt-3.5-turbo with a system role matching that of the human writers, and to be of comparable length ($\sim$500 words). The resulting dataset, which we name GPT-WritingPrompts, is a useful source of free-form natural language short stories with matching human and machine responses to the same prompt.
@@
    Analyzing the narrative characteristics of stories written by both sets of writers along six dimensions. The first three are the affective dimensions of valence (positive–negative), arousal (active–passive), and dominance (powerful–weak), which capture the three fundamental dimensions of connotative meaning as formulated by Osgood et al. (1958). We also look at three other axes of character portrayal that are often noted in prior work for stereotypical connotations along gender lines: intellect, appearance, and power.
@@
    Quantifying differences along these dimensions, in both human-written and GPT-generated stories, when grouped by the narrative voice and the gender of the main protagonist of the story. Stories are categorized as being in first-person, second-person, or third-person, based on the dominant pronouns used in them. Third-person stories are further categorized as featuring a male or female protagonist (or Other, if pronouns do not fall exclusively into one of these categories).
  4. 4.
    Comparing the consistency of biases in portrayal when analyzed using different methods proposed in past literature. We unify multiple methods from past work to (a) identify protagonist attributes given a story, and (b) score a set of lexical attributes along a dimension of interest (such as valence or power), in a single pipeline, enabling us to study the sensitivity of our measures of bias to the methods that are used to quantify them.
Our results show that GPT-generated stories differ significantly from human-written ones along all six dimensions, when compared at both the aggregate-level (Sections 5.2) and at the instance-level (Section 5.3). On average, GPT-generated stories use more words associated with a positive valence and higher dominance compared to human stories, but lower arousal (less activity), and fewer words that relate to appearance and intellect on average. Both sets of stories, however, display similar biases towards their male and female protagonists: female protagonists are more positive, less aroused, less in control, have a higher proportion of appearance-related words associated with them, and a lower proportion of intellect-related words compared to male protagonists. These results support prior findings on how neural models of language generation, both for stories and otherwise, differ from those of humans Lucy and Bamman (2021); Huang et al. (2021); Giulianelli et al. (2023).
@@ 2.1 Generative Models of Natural Language
Text generation, both conditional and open-ended, has been a core area of NLP research for many decades. The development of neural transformer models, pretrained with the language modeling objective on large amounts of textual data, has led to a exponential improvement in the fluency, coherence, and versatility of machine-generated text Li et al. (2022). Popular models include the GPT family of models from OpenAI [^2] https://platform.openai.com/docs/models, Llama models from Meta AI Touvron et al. (2023), and the Claude family of models from Anthropic AI Anthropic (2024), among others. These models are able to generate long-form text (up to thousands of tokens in length) in response to a user-provided input prompt, which has led to their adaptation as the base models for several NLP tasks, including translation, question answering, summarization, and controlled text generation Brown et al. (2020).
Parallel to these advances are works that study the probabilistic algorithms behind the generative processes, and how they differ from that of human language generation. Generally, given a context, this involves designing metrics that compare how similar machine generations are to a set of reference human generations (reference-based evaluation). Metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), MoverScore Zhao et al. (2019), and BertScore Zhang et al. (2020) fall under this umbrella; however, for open-ended tasks like story generation, they do not capture the full range of permissible outputs. Reference-free evaluations, on the other hand, train models to discriminate between human and machine generations Hashimoto et al. (2019); De Mattei et al. (2021). In recent times, LLMs themselves have been proposed as evaluators of the quality of their generations Zhong et al. (2022); Eldan and Li (2023).
A third approach, closest to what we follow in this work, is studying the statistical distributions of various text features (type-to-token ratio, word and sentence lengths) in the generations of the two processes (humans and machine). These measures can be computed and compared both at the corpus-level, and at the instance-level Meister and Cotterell (2021); Pillutla et al. (2021); Giulianelli et al. (2023).
Another equally important area of research is to be quantify the default generative behaviours of these models, understand the biases exhibited by them towards or against particular cultures, ideologies, or products, and whether they reinforce harmful stereotypes about certain groups, particularly those that have historically been marginalized. This is the primary motivation behind our work, which seeks to quantify how certain aspects of storytelling differ between human writers and machine generations.
@@ 2.2 Biases in Stories
Research into biases in stories, both human and artificial, has focused largely on differences in the portrayal of characters belonging to different demographic groups. Studies on the portrayal of gendered groups in particular, though mostly limited to the male/female binary, have demonstrated that female characters tend to be portrayed as being more emotional, having less agency and power, and more prone to being described by physical attributes. Fast et al. (2016b) create lexicons of words associated with several categories of behavioural stereotypes, such as violent, intelligent, domestic, emotional, etc. Linguistic dependency relations are then used to measure the level of association of verbs and adjectives associated with character to each of these categories.
Sap et al. (2017) create lexicons of verbs that quantify the distribution of power and agency between the subject (agent) and object (theme) of these verbs. They use these lexicons to measure the power and agency assigned to characters in film scripts via their dialogue, and find that male characters are generally given more agency.
Lucy and Bamman (2021) also use linguistic dependencies to quantify the level of association of gendered character groups with words related to appearance, intellect, and power, for both human-written and GPT-3 generated story excerpts. They find that generated characters have lower power and intellect scores compared to human stories, but both sets of stories display similar gender differences – female characters are associated with more appearance-related words, and male characters with higher levels of power.
Huang et al. (2021) use commonsense inference with COMET to derive more implicit indicators of bias that are not explicitly stated, and therefore will be missed by methods that use linguistic dependencies. They analyze implicit bias with the ROCStories dataset, which consists of short 5-sentence narrative stories, for both human and GPT-2 generations. Along with appearance, intellect, and power, they study the two affective dimensions of valence and arousal. They also find that female characters have higher associations with appearance, lower associations with intellect, and lower intensities of power; however, they do not see significant differences in valence and arousal.
Here, we follow prior work in analyzing appearance, intelligence, and power as aspects of character portrayal. We further include the three affective dimensions of valence, arousal, and dominance (dominance is closely related to power). We compare and contrast the results of characterizing a protagonist by both explicit linguistic markers (verbs and adjectives), as well as commonsense inferences with COMET.
@@ 3 The GPT-WritingPrompts Dataset
Fan et al. (2018) collect a large dataset of 300K stories written in response to 97,223 writing prompts from the Reddit r/WritingPrompts forum, henceforth referred to as the WritingPrompts dataset. The forum is a place for writers to gather inspiration for stories – users post short, 1-2 sentence prompts designed to invoke a certain scenario or set the premise for a potentially-exciting narrative to unfold, and other forum participants respond with short stories relating to, or following, the premise set-up in the prompt.
@@
We collect all prompts and associated stories in the training set of the WritingPrompts dataset. Each prompt is associated on average with 2.8 stories, and the stories average 675 tokens in length (Table 1).
@@ 3.1 Generating Artificial Stories
The generation interface for GPT models allows us to specify a system role that helps set the behaviour of the generative model. These roles influence the “persona" assumed by the model as it generates text for subsequent input contexts – one can specify a certain style of talking, or a level of intelligence, or formatting specifications for the output. Considering that our domain of narratives is quite specific – short stories written on an internet forum – we test two variations of the system role:
@@
With the system role specified, we obtain generations for each story prompt by providing the following text as context to the model: “Write a 500 word story for the following prompt: <text of the prompt>". We obtain an average of 2 stories for each input prompt, with a softmax temperature of 0.95 for high variability. An initial, qualitative examination of generated stories did not indicate any obvious stylistic differences between the two system roles.
We term the combined dataset of writing prompts paired with human-written and machine-generated short stories the GPT-WritingPrompts dataset. It allows us to systematically examine differences in the responses of human writers and a large language model to the same input contexts, in a narrative, length-constrained creative writing task.
Table 1 reports some basic statistics of the human-written and GPT-generated subsets of our dataset. Note that despite the specified 400-word limit, the model tends to generate slightly longer stories, averaging at  530 words.
@@ 4 Methodology
We now describe the analysis pipeline for stories, which is applied to the outputs of both groups of writers, human and GPT-3.5.
Given an input story and a dimension of interest – say, valence – our aim is to obtain a numerical score that represents the intensity, or level, of that dimension ascribed to the protagonist of the story. A story with a very positive main character will therefore be represented by a high valence score, and one where the protagonist is associated with negative sentiment terms will have a low valence score. We break this process down into a series of computational steps.
@@ 4.1 Characterizing the Point-of-View
Character roles in literature can often be complex and gray; here, we opt for the simpler strategy of declaring the most frequently mentioned character entity as the protagonist. We characterize the point-of-view (PoV) of the story by identifying the grammatical person of narrative voice (first, second, or third person), and, where possible, identifying the referential gender of third-person protagonists.
@@
We then retrieve the pronominal mentions of the protagonist clusters and compare them to a predefined list of pronominal terms associated with each point-of-view category (the full set of terms is listed in Table 5 of the Appendix). We first separate out stories in the first person (FP) and second person (SP) by identifying cases where the associated (FP and SP) pronouns account for more than half of all pronominal mentions of the protagonist.
Stories that don’t fall into the above two categories are considered to be in the third person (TP). Here, we label a story’s protagonist as Male (TP-M) if their referent pronouns are exclusively in the set of Male-associated pronouns, and Female if they are exclusively referred to by Female-associated pronouns (TP-F). If the pronouns do not fit into either category, then the protagonist is categorized under a third group, labelled Other. We observed that this often results from mentions such as “It", “They", and “Us", or from clusters containing pronominal terms from multiple pronominal categories.
@@ 4.2 Extracting Protagonist Attributes
We now proceed to extract a list of attributes that describe the protagonist and their actions in the narrative. We identify two main methods to obtain these tokens from the text, inspired by prior work:
Linguistic Dependencies with spaCy:
We use the linguistic dependency relations among the tokens of each sentence in the story to identify the verbs and adjectives attached to each identified protagonist mention. This corresponds to the method used in Lucy and Bamman (2021).
Implicit Attributes with COMeT:
We follow the commonsense reasoning-based method described in Huang et al. (2021) to infer attributes beyond lexical indicators from the story text. COMeT is a generative knowledge base completion model that, given a sentence containing one or more entities, generates short phrases that describe the underlying relations between the entities. The motivation for this method is to be able to identify underspecified and implicit indicators of bias.
We present sample extracts from three stories from our dataset, along with the attributes obtained by each of these pipelines, in Table 6 of the Appendix.
@@ 4.3 Dimensions of Entity Portrayal
One can imagine several abstract dimensions that might be of interest when characterizing a protagonist: whether they are good or bad, powerful or weak, happy or sad, superficial or authentic, and so on.
Several seminal works on the semantics of words Osgood et al. (1958); Russell (1980) have formulated a view of the connotative meanings of words (contrasted with the denotative meaning, used to refer to the literal word meaning as found in a dictionary) as being composed of three primary dimensions: valence (the evaluative good–bad axis), arousal (also called activation, an indication of active–passive), and dominance (potency, characterized along the strong–weak axis). We take these three dimensions as our first set of axes along which to quantify character portrayal.
A primary lexical resource assessing word meaning along these dimensions is the NRC-VAD lexicon, which consists of a list of $\sim$20,000 common English words associated with real-valued scores for valence, arousal, and dominance. The scores fall in the range between 0 (indicating least valence/highest negativity, lowest arousal/highest passivity, and lowest dominance/least in-control) and 1 (most positive, active, and in-control).
Apart from this, we consider two binary quantifiers of character portrayal, appearance and intellect. These two dimensions, along with power, have largely been the focus of prior work in this domain Lucy and Bamman (2021); Huang et al. (2021). [^3] We note that the dimension of power is conceptually very close to dominance, and we expect to be capturing similar aspects of character portrayal with these two dimensions. In our experiments, we consider power as a sixth pseudo-dimension of analysis to stay consistent with prior work, and also use it to test the consistency of our conclusions with two different lexicons for aspect of power/dominance.
We use similar lexical resources to prior works in order to quantify appearance and intellect: a list of terms related to the concepts beautiful and sexual to characterize appearance (sourced from the Empath lexical resource from Fast et al. (2016a)), and a list of terms related to the intellect category from Fast et al. (2016b). Unlike the VAD lexicons, terms in these lexicons are not scored along a real-values bipolar scale; rather, they function as categorical, boolean indicators of association with the dimension, whether in a negative light ("the protagonist was ugly") or a positive light ("the protagonist was smart").
Power, on the other hand, is a boolean, bipolar dimension represented by a list of terms that represent the concepts powerful and weak, sourced from the lexicon from Fast et al. (2016b).
We list some representative terms along with their scores for each dimension in Table 7 of the Appendix.

### v2

@@ Abstract
The improved generative capabilities of large language models have made them a powerful tool for creative writing and storytelling. It is therefore important to quantitatively understand the nature of generated stories, and how they differ from human storytelling. We create the GPT-WritingPrompts dataset, which pairs short stories written by Reddit users in response to 97,219 prompts with comparable generations from GPT-3.5. We quantify the emotional and descriptive features of storytelling for both generative processes, human and machine, along a set of six dimensions. We find that generated stories differ significantly from human stories along all six dimensions, and that human and machine generations display similar biases when grouped according to the narrative point-of-view and gender of the main protagonist. We make our dataset and code publicly available [^1] https://github.com/KristinHuangg/gpt-writing-prompts.
@@ 1 Introduction
Storytelling is fundamentally human, and a rich representation of the beliefs, practices, and languages of communities. The domain of literature is therefore a source of cultural information for researchers in the social sciences, from historians to linguists Piper et al. (2021); Ziems et al. (2024). The characters [^2] We use the term ‘character’ in this work to refer to fictional people, i.e, named entities that speak and act in stories. in a story are often central to this process: protagonists are usually depicted with socially and personally desirable traits like kindness, intelligence, and strength, and villains with undesirable characteristics (greed, cruelty, apathy, weakness). Consequently, we have seen extensive research on NLP methods that can model and understand these aspects of stories, and draw inferences about social biases represented in them Fast et al. (2016b); Lucy and Bamman (2021); Sap et al. (2022).
Large language models (LLMs), pretrained on billions of tokens of text from the internet, are now able to generate several thousand words of coherent output in response to user-specified input contexts Bommasani et al. (2021). Naturally, these systems have become popular tools for a number of tasks involving text generation, both controlled and open-ended. e.g., dialogue systems, code generation, summarization, question answering, and creative writing. However, neural models have been shown to replicate and amplify the biases encoded in their training data Bender et al. (2021). It is therefore interesting – and important – to understand the nature of stories that are told by these models, and how they differ from human models of storytelling Chakrabarty et al. (2023); Xie et al. (2023); Chakrabarty et al. (2024); Breithaupt et al. (2024).
@@
    Extending the Reddit WritingPrompts dataset Fan et al. (2018) to include generations from the OpenAI GPT-3.5 model (gpt-3.5-turbo). The original dataset consists of 97,222 unique writing prompts, and multiple stories written by Reddit users in response to each prompt. We generate a corresponding set of artificial stories by prompting gpt-3.5-turbo with a role matching that of the human writers, and for a comparable length ($\sim$500 words). The resulting dataset, GPT-WritingPrompts, is a useful source of free-form natural language short stories with matching human and machine responses to the same prompt.
@@
    Analyzing the narrative characteristics of stories written by both sets of writers along six dimensions. The first three are the affective dimensions of valence (positive–negative), arousal (active–passive), and dominance (powerful–weak), which capture the three fundamental dimensions of connotative meaning as formulated by Osgood (1957). We also look at three other axes of character portrayal that are often noted in prior work for stereotypical connotations along gender lines: intellect, appearance, and power.
@@
    Quantifying biases in character portrayal along these dimensions, in both human-written and GPT-generated stories, when grouped by the narrative voice and the gender of the main protagonist of the story. Stories are categorized as being in first-person, second-person, or third-person, based on the dominant pronouns used in them. Third-person stories are further categorized as featuring a male or female protagonist (or Other, if pronouns do not fall exclusively into one of these categories).
We show that GPT-generated stories differ significantly from human-written ones along all six dimensions, when compared at both the aggregate-level (Sections 5.2) and at the instance-level (Section 5.3). On average, GPT-generated stories use more words with a positive valence and higher dominance compared to human stories, but lower arousal (less activity), and use fewer words that relate to appearance and intellect. Both sets of stories, however, display similar biases towards their male and female protagonists: female protagonists are more positive, less aroused, less in control, are associated with a higher proportion of appearance-related words, and a lower proportion of intellect-related words compared to male protagonists. This supports prior findings on how neural models of language generation differ from those of humans Lucy and Bamman (2021); Huang et al. (2021); Giulianelli et al. (2023); Breithaupt et al. (2024).
@@ 2.1 Generative Models of Natural Language
Text generation, both conditional and open-ended, has been a core area of NLP research for many decades. The development of neural transformer models, pretrained with the language modeling objective on large amounts of textual data, has led to exponential improvements in the fluency, coherence, and versatility of machine-generated text Li et al. (2024). Popular models include the GPT family of models from OpenAI [^3] https://platform.openai.com/docs/models, Llama models from Meta AI Touvron et al. (2023), and the Claude family of models from Anthropic AI Anthropic (2024), among others. These models are able to generate long-form text (up to thousands of tokens in length) in response to a user-provided input prompt, which has led to their adaptation as the base models for several NLP tasks, including translation, question answering, summarization, and controlled text generation Brown et al. (2020).
Parallel to these advances are studies of the probabilistic algorithms behind generative processes of these models, and how they differ from human language generation. Given a context, this involves designing metrics that compare how similar machine generation is to a set of reference human generations (reference-based evaluation). Another approach, closer to ours and more appropriate for open-ended text generation, is to study the statistical distributions of various text features (type-to-token ratio, word and sentence lengths) in the two generation processes (humans and machine). These measures can be computed and compared both at the corpus-level, and at the instance-level Meister and Cotterell (2021); Pillutla et al. (2021); Giulianelli et al. (2023).
Another equally important area of research is to quantify the default generative behaviours of these models, understand the biases exhibited by them towards or against particular cultures, ideologies, or products, and whether they reinforce harmful stereotypes about certain groups, particularly those that have historically been marginalized.
@@ 2.2 Biases in Stories
Research into biases in stories, both human and artificial, has focused largely on differences in the portrayal of characters belonging to different demographic groups. Studies on the portrayal of gendered groups in particular, though mostly limited to the male/female binary, have demonstrated that female characters tend to be portrayed as being more emotional Fast et al. (2016b); Zhou et al. (2022), having less agency and power Sap et al. (2017), and are more prone to being described by physical attributes Lucy and Bamman (2021); Huang et al. (2021); Begus (2023); Zhao et al. (2024).
A common strategy for measuring such biases is to first extract descriptive lexical attributes associated with gendered characters from the story text, and then quantify the bias of these terms towards a certain lexical dimension, such as power or appearance, using an associated lexicon. Fast et al. (2016b) create lexicons of words associated with several categories of behavioural stereotypes, such as violent, intelligent, domestic, emotional, etc. Sap et al. (2017) create lexicons of verbs that quantify the distribution of power and agency between the subject (agent) and object (theme) of verbs. Emotion lexicons have been created for many emotion dimensions, and in multiple languages, such as LIWC Tausczik and Pennebaker (2010), WordNet-Affect Bobicev et al. (2010), SentiWordNet Baccianella et al. (2010), VADER Hutto and Gilbert (2014), and the NRC suite of emotion and affect lexicons Mohammad (2018).
Here, we analyze appearance, intelligence, and power as aspects of character portrayal. We also include the three affective dimensions of valence, arousal, and dominance. We compare and contrast the results of characterizing a protagonist by explicit linguistic markers (verbs and adjectives), as well as commonsense inferences with COMET.
@@ 3 The GPT-WritingPrompts Dataset
Fan et al. (2018) collected 300K stories written in response to 97,223 prompts from the Reddit r/WritingPrompts forum, referred to as the WritingPrompts dataset. Users of the forum post short prompts designed to invoke a scenario or set the premise for a narrative to unfold, and other forum participants respond with short stories relating to, or following, the premise set-up in the prompt.
@@
We collect all prompts and associated stories in the training set of WritingPrompts. Each prompt is associated on average with 2.8 stories, and the stories average 675 tokens in length (Table 1).
@@ 3.1 Generating Artificial Stories
GPT models allow us to specify a system role that helps set the behaviour of the generative model. These roles influence the “persona" assumed by the model as it generates text for subsequent input contexts – one can specify a certain style of talking, or a level of intelligence, or formatting specifications for the output. Considering that our domain of narratives is quite specific – short stories written on an internet forum – we test two variations of the system role:
@@
An initial, qualitative examination of generated stories for a random sample of 100 prompts did not indicate any obvious stylistic differences between the two system roles; we therefore stick with the Redditor role for the rest of the generations.
With the system role specified, we obtain generations for each story prompt by providing the following text as context to the model: “Write a 500 word story for the following prompt: <text of the prompt>". We obtain an average of 2 stories for each input prompt, with a softmax temperature of 0.95 for high variability.
We term the combined dataset of writing prompts paired with human-written and machine-generated short stories the GPT-WritingPrompts dataset. It allows us to systematically examine differences in the responses of human writers and an LLM to the same input contexts, in a narrative, length-constrained creative writing task.
Table 1 reports some basic statistics of the human-written and GPT-generated subsets of our dataset. Note that despite the specified 500-word limit, the model tends to generate slightly longer stories, averaging at  530 words.
@@ 4 Methodology
Given an input story and a lexical dimension of interest – say, valence – our aim is to obtain a numerical score that represents the intensity, or level, of that dimension ascribed to the protagonist of the story. A story with a very positive main character will therefore be represented by a high valence score, and one where the protagonist is associated with negative sentiment terms will have a low valence score. We break this process down into a series of computational steps.
@@ 4.1 Characterizing the Point-of-View
Character roles in literature can often be complex and gray; here, we opt for the simpler strategy of declaring the most frequently mentioned character entity as the protagonist. We characterize the point-of-view (PoV) of the story by identifying the grammatical person of the narrative voice (first, second, or third person), and, where possible, identifying the referential gender of third-person protagonists.
@@
We retrieve the pronominal mentions of the protagonist clusters and compare them to a predefined list of pronominal terms associated with each point-of-view category (the full set is in Table 5 of the Appendix).
We first separate out stories in the first person (FP) and second person (SP) by identifying cases where the associated (FP and SP) pronouns account for more than half of all pronominal mentions of the protagonist. Stories that don’t fall into the above two categories are considered to be in the third person (TP). Here, we label a story’s protagonist as Male (TP-M) if their referent pronouns are exclusively in the set of Male-associated pronouns, and Female if they are exclusively referred to by Female-associated pronouns (TP-F). If the pronouns do not fit into either category, then the protagonist is categorized under a third group, labelled Other. We observed that this often results from mentions such as “It", “They", and “Us", or from clusters containing pronominal terms from multiple pronominal categories.
@@ 4.2 Extracting Protagonist Attributes
We extract a list of attributes from each story that describe the protagonist and their actions. We use two methods to obtain these tokens from the text:
Linguistic Dependencies (with spaCy) among the tokens of each sentence to identify the verbs and adjectives attached to each protagonist mention. This corresponds to the method used by Lucy and Bamman (2021).
Implicit Attributes (with COMeT). We follow the commonsense reasoning-based method described by Huang et al. (2021) to infer attributes beyond lexical indicators from the text. COMeT is a generative knowledge base completion model that, given a sentence containing one or more entities, generates short phrases that describe the underlying relations between those entities. The motivation is to be able to identify underspecified and implicit indicators of bias.
We present sample extracts from three stories from our dataset, along with the attributes obtained by both methods, in the Appendix (Table 6).
@@ 4.3 Dimensions of Entity Portrayal
One can imagine several abstract dimensions that might be of interest when characterizing a protagonist: whether they are good or bad, powerful or weak, happy or sad, superficial or authentic, etc.
Several seminal works on the semantics of words Osgood (1957); Russell (1980) have formulated a view of the connotative meanings of words (contrasted with the denotative meaning, used to refer to the literal word meaning as found in a dictionary) as being composed of three primary dimensions: valence (the evaluative good–bad axis), arousal (also called activation, an indication of active–passive), and dominance (potency, characterized along the strong–weak axis). We take these three dimensions as our first set of axes along which to quantify character portrayal.
The NRC-VAD lexicon Mohammad (2018) consists of $\sim$20,000 common English words associated with real-valued scores for valence, arousal, and dominance. The scores fall between 0 (indicating least valence/highest negativity, lowest arousal/highest passivity, and lowest dominance/least in-control) to 1 (most positive, active, and in-control).
We also consider two binary quantifiers of character portrayal, appearance and intellect. These two dimensions, along with power, have largely been the focus of prior work Lucy and Bamman (2021); Huang et al. (2021). [^4] The dimension of power is conceptually close to dominance, and we expect to be capturing similar aspects of character portrayal with these two dimensions. Here, we consider power as a sixth pseudo-dimension of analysis to stay consistent with prior work, and also use it to test the consistency of our conclusions with two different lexicons for the aspect of power/dominance. We use similar lexical resources to prior works in order to quantify appearance and intellect: a list of terms related to the concepts beautiful and sexual to characterize appearance (sourced from the Empath lexicon from Fast et al. (2016a)), and a list of terms related to the intellect category from Fast et al. (2016b). Unlike the VAD lexicons, terms in these lexicons are not scored along a real-valued bipolar (low valence vs high valence) scale; rather, they function as categorical, unipolar, boolean indicators of association with the dimension, whether in a negative light ("the protagonist was ugly") or a positive light ("the protagonist was smart"). Power, on the other hand, is a boolean, bipolar dimension represented by a list of terms that represent the concepts powerful (high power) and weak (low power), sourced from the lexicon of Fast et al. (2016b).
Representative terms and their scores for each dimension are in Table 7 of the Appendix.